OpenStudy (anonymous):

Express the integrand as a sum of partial fractions and evaluate the integral. Integral from 0 to 1 of (x^3)/ (x^2+8x+16) dx

OpenStudy (anonymous):

\[\int\limits_{0}^{1} \frac{ x^{3} }{ x ^{2}+8x+16}dx\] I've gotten as far as : \[\int\limits_{0}^{1} x+1+\frac{ 256 }{ (x+4)^{2} } +\frac{ 51 }{ (x+4)} dx\]

OpenStudy (solomonzelman):

I think that is incorrect...

OpenStudy (anonymous):

Ok, then how do I figure out A and B if the denominator is a perfect square?

OpenStudy (solomonzelman):

before I do division, this is how you do it one fraction is (x+4) and the other is (x+4)^2

OpenStudy (solomonzelman):

(I am talking denominators)

OpenStudy (anonymous):

Right

OpenStudy (solomonzelman):